### v1

@@ 2.1 Supervised Desmoking
Several studies [58, 75, 55, 59, 40, 71, 60] have endeavored to address the surgical smoke removal problem. Traditional approaches [75, 55, 58] estimate the transmission components based on the atmospheric scattering model, but easily lead to color and structure artifacts. For example, Lin et al. [36] and Ma et al. [40] modify U-net [48] architecture for surgical desmoking. Wang et al. [59] propose an encoder-decoder architecture with a Laplacian image pyramid decomposition strategy. Wang et al. [60] employ Swin Transformer blocks [35] to enhance feature extraction. Zheng et al. [71] further develop a system jointly detecting and removing surgical smoke. However, these methods generally simulate surgery smoke for training. Due to the domain gap between the synthetic smoke and the real-world one, they can’t generalize well to real surgery scenes.
@@ 2.2 Supervised Dehazing
Dehazing [5, 29, 19, 11, 63, 30, 47, 67, 37, 64, 72, 45, 20] is of great relevance to surgical desmoking, which aims to recover clean components from hazy ones affected by adverse weather conditions. Li et al. [29] reformulate the atmospheric scattering model and propose AODNet for end-to-end image dehazing. UHD [63] introduces the infinite approximation of Taylor’s theorem with Laplacian pyramid pattern to handle hazy images. DeHamer [19] further combine CNN [48] and vison Transformer [13] together for performance improvement. Compared to the above single-image methods, video ones [30, 47, 67, 67, 37, 64] can leverage temporal clues between consecutive frames for more effective restoration. Li et al. [30] first build an united video detection and dehazing framework that focuses on temporal information fusion. Ren et al. [47] incorporate global semantic priors for smooth transmission map estimation. Recently, MAPNet [64] explores physical haze priors to guide spatial information extraction and proposes a spatial-temporal alignment strategy to guide temporal features aggregation, achieving state-of-the-art performance. However, employing their pre-trained models directly does not produce satisfactory desmoking results due to the discrepancy between atmospheric haze and surgical smoke. Moreover, the lack of real-world paired data severely restricts the possibility of starting training from scratch. In this work, we propose a self-supervised framework for video desmoking (SelfSVD) to address these issues.
@@ 2.3 Unsupervised Desmoking and Dehazing
In contrast to supervised methods, unsupervised ones [73, 57, 24, 43, 50, 8, 53, 31, 32, 70, 65, 33, 15, 66, 16, 9] can be trained without paired supervision. For surgical smoke removal, Cyclic-DesmokeGAN [57] proposes a real-world single-image desmoking model based on CycleGAN [74]. Desmoke-LAP [43] and DCP-Pixel2Pixel [50] introduce dark channel prior [21] into loss and network design, respectively. MS-CycleGAN [53] proposes a multi-stage method that adapts a model pre-trained on synthetic data to real-world ones.
@@ 3.1 Taking PS Frame as Supervision
Given a laparoscopic smoky video consisting of $N$ frames $\{\mathbf{S}_{i}\}^{N}_{i=1}$, video desmoking aims to restore the corresponding clean components ${\{\mathbf{\hat{I}}_{i}\}}^{N}_{i=1}$ with temporal clues, which can be written as,
@@
$\mathcal{L}_{rec}=\sum_{i=1}^{N}||\mathbf{V}_{i}\odot(\mathbf{\hat{I}}_{i\rightarrow ps}-\mathbf{S}_{ps})||_{1},$ (5)
where $\mathbf{V}_{i}$ is a mask that indicates the valid positions of optical flow, and $\odot$ is a pixel-wise multiply operation.
@@ 3.2 Taking PS Frame as Reference Input
Masking Strategy. In regions with significantly inaccurate optical flow, we prefer to forget the reference information to avoid output contents being inconsistent with smoky inputs. As shown in Fig. 4, we suggest generating a mask $\mathbf{M}_{i}$ to indicate these regions in $\mathbf{F}_{ref\rightarrow i}$, and then exclude the corresponding features, which can be written as,
@@
In particular, a mask generator is deployed to get $\mathbf{M}_{i}$. First, we align $\mathbf{S}_{ref}$ to $\mathbf{S}_{i}$, obtaining the warped reference image $\mathbf{S}_{ref\rightarrow i}$. Then, we process $\mathbf{S}_{ref\rightarrow i}$ and $\mathbf{S}_{i}$ with DCP [21] and large-kernel Gaussian blur operations to alleviate the disturbance of surgery smoke. Next, we divide $\mathbf{S}_{ref\rightarrow i}$ and $\mathbf{S}_{i}$ into $P$ non-overlapping patches respectively, i.e., $\{{\mathbf{S}_{ref\rightarrow i}^{p}\}}^{P}_{p=1}$ and $\{{\mathbf{S}_{i}^{p}\}}^{P}_{p=1}$. For each patch pair $\mathbf{S}_{ref\rightarrow i}^{p}$ and $\mathbf{S}_{i}^{p}$, their structure should be significantly distinct when optical flow is estimated inaccurately. So we adopt structural similarity (SSIM) metric [61] to detect such areas, which can be written as,
@@
$m_{i}^{p}$ is the value of $\mathbf{M}_{i}$ in the $p$-th patch. $\max$ and $\operatorname{sgn}$ denote the maximum and sign function, respectively. $\epsilon$ is a threshold and is set to 0.92. The patch size is set to $8$. Please see the mask visualizations and more details in the Suppl.
Regularization Strategy. The areas where optical flow is slightly wrong may lead to potential trivial solutions, as they are not easily detected explicitly. We use $\ell_{1}$ regularization [12, 3] from conventional wisdom in machine learning to inhibit the effects of these areas. The regularization loss $\mathcal{L}_{reg}$ can be written as,
$\mathcal{L}_{reg}=||\tilde{\mathbf{F}}_{ref\rightarrow i}||_{1}.$ (9)
@@ 3.3 Enhancing PS Frame as Supervision
Figure 5: The illustration of enhancing PS frame ($\mathbf{S}_{ps}$) as supervision. We regard $\mathbf{S}_{ps}$ as a frame with less smoke and feed it into a pre-trained SelfSVD model, generating a cleaner result $\mathbf{S}_{ps}^{\ast}$. Then, $\mathbf{S}_{ps}^{\ast}$ is taken as improved supervision to fine-tune the SelfSVD model by $\mathcal{L}_{rec}$ and $\mathcal{L}_{GAN}$ (i.e., replacing $\mathbf{S}_{ps}$ with $\mathbf{S}_{ps}^{\ast}$ in Eq. 5 and Eq. 10), getting an improved model named ${}^{\ast}$.
Taking this into account, we improve the training of SelfSVD by enhancing $\mathbf{S}_{ps}$ as better supervision, dubbed ${}^{\ast}$, as shown in Fig. 5. In particular, a SelfSVD model is first pre-trained under the supervision of $\mathbf{S}_{ps}$. Then, we regard $\mathbf{S}_{ps}$ as a frame with less smoke and feed it into the pre-trained SelfSVD, getting a cleaner result $\mathbf{S}_{ps}^{\ast}$. Finally, a ${}^{\ast}$ model is obtained by fine-tuning the pre-trained SelfSVD with $\mathbf{S}_{ps}^{\ast}$ as supervision, and the original SelfSVD model can be detached during testing.
@@ 3.4 Network Architecture and Learning Objective
where $\lambda_{reg}$ and $\lambda_{GAN}$ are set to 0.05 and 1.0, respectively. When training $*{}^{*}$ model, $\mathbf{S}_{ps}$ in Eq. 12 is replaced with $\mathbf{S}_{ps}^{*}$, as stated in Sec. 3.3.
@@
(Ours) ${}^{\ast}$ 24.58 0.6279 0.4193 4.72 3.86 996 0.18 15.58
4 LSVD Dataset
5 Experiments
5.1 Implementation Details
For optical flow estimation, we adopt a pre-trained PWC-Net [54]. During training, we randomly crop patches and augment them with random flips. The batch size is set to 4 and the patch size is set to $256\times 256$. SelfSVD is trained with ADAM optimizer  [28] with ${\beta}_{1}=0.9$ and ${\beta}_{2}=0.99$ for 100k iterations. Cosine annealing strategy  [38] is employed to decrease the learning rate from $1\times 10^{-4}$ to $1\times 10^{-7}$ steadily. For the training of ${}^{\ast}$, we fine-tune the pre-trained SelfSVD model for additional 40k iterations and set the initial learning rate to $5\times 10^{-5}$. All experiments are conducted with PyTorch [44] on a single Nvidia GeForce RTX A6000 GPU.
SelfSVD and ${}^{\ast}$ keep the same number of residual blocks [22] as BasicVSR++ [7], and the computation costs are generally similar to the video processing methods. Moreover, to make the model inference cost consistent with some single-image processing methods, we present the lightweight versions by reducing the numbers and channels of residual blocks, dubbed SelfSVD-S and ${}^{\ast}$-S respectively. Please see more details in the Suppl.
[…]
Figure 7: Qualitative comparison on LSVD dataset. Our methods remove more clean smoke and recover more realistic details.
@@ 5.3 Comparison with State-of-the-Art Methods
Related learning-based methods are either trained on synthetic data in a supervised manner[29, 59, 11, 63, 19, 20, 6, 7, 64], or trained in a unpaired manner[50, 52, 70, 43]. For fair comparisons, we retrain their models on our LSVD dataset. For training supervised methods, there is no paired ground-truth. Instead, we take the aligned PS frame as the supervision, which is the same as the setting of training our models as stated in Sec. 3.1. Thus, they become self-supervised methods. For unpaired learning methods, we take the original PS frame as their unpaired supervision. Overall, we compare our methods (i.e., SelfSVD-S and ${}^{\ast}$-S, SelfSVD, ${}^{\ast}$) against 16 related state-of-the-art ones, including 2 unsupervised image processing methods (i.e., PSD [9], DCP [21]), 4 unpaired image ones (i.e., DCP-Pixel2Pixel [50], DistentGAN [52], and Desmoke-LAP [43], RefineNet [70]), 7 self-supervised image ones (i.e., UHD [63], MSDesmoking [59], Wang et al. [60] MSBDN [11], AODNet [29], DADFNet [20], Dehamer [19],) and 3 self-supervised video ones (i.e., BasicVSR [6], BasicVSR++ [7], and MAPNet [64]). We do not compare with unsupervised and unpaired video ones, as few works explore that as far as we know. Beyond conducting experiments on our dataset, we provide more comparisons on a synthetic dataset in the Suppl.
Quantitative Comparison. Tab. 1 summarizes the quantitative comparison results on LSVD dataset. First, video processing methods generally perform better than single-image ones, which demonstrates the benefits of utilizing temporal clues for surgical smoke removal. Second, our methods outperform state-of-the-art ones by a large margin. Among all competing methods, BasicVSR++ [7] has achieved the best PSNR score of 23.35dB. Our SelfSVD and ${}^{\ast}$ achieve PSNR gains of 0.88dB and 1.23dB over BasicVSR++ [7], respectively. Although PSD [70] and DCP [21] get better FADE scores than ours, they generally over-process smoky videos and result in over-saturated colors (as shown in Figs. 6 and 7), leading to poor NIQE and PI scores, as well as unsatisfactory visual effects. Third, with lower computation cost and fewer model parameters, the proposed SelfSVD-S and ${}^{\ast}$-S still perform well, which further illustrates the effectiveness of our methods.
Qualitative Comparison. Due to space limitation, we only provide qualitative results of some methods with better quantitative scores in Figs. 6 and 7. It can be seen that the compared methods often introduce color distortion, over-smoothing, or smoke preservation in their results. Instead, our methods remove more smoke and restore more details that are consistent with PS frames. More qualitative comparison results can be seen in the Suppl.
@@ 6.1 Effect of Taking PS Frame as Supervision
6.2 Effect of Taking PS Frame as Reference Input
Note that PS frame is available during training and testing. We take it as an additional reference (Ref) input. We perform experiments to evaluate its effectiveness by removing the Ref. The quantitative results are shown in Tab. 3 and the qualitative ones are shown in Fig. 8. It can be seen that the additional Ref enables 0.67dB PSNR improvements. Besides, Fig. 8 shows that it leads to cleaner smoke removal and finer-scale detail recovery, especially in areas with dense smoke.
Additionally, we also conduct experiments to validate whether PS frame is a suitable choice to be used as Ref. For comparison, we replace Ref with the first smoky frame, as it generally has less smoke than subsequent smoky ones. As shown in Tab. 3, the replacement leads to 0.28dB PSNR drop, which illustrates the frames before the activations of high-energy devices are preferable to be taken as Ref.

### v2

@@ 2.1 Supervised Desmoking
Several studies [58, 75, 55, 59, 40, 71, 60] have endeavored to address the surgical smoke removal problem. Traditional approaches [75, 55, 58] estimate the transmission components based on the atmospheric scattering model, but easily lead to color and structure artifacts. Later works [59, 40, 36, 71, 60] adopt learn-based manners and design various networks based on convolutional neural networks (CNN) and vision transformers [13]. For example, Lin et al. [36] and Ma et al. [40] modify U-net [48] architecture for surgical desmoking. Wang et al. [59] propose an encoder-decoder architecture with a Laplacian image pyramid decomposition strategy. Wang et al. [60] employ Swin Transformer blocks [35] to enhance feature extraction. Zheng et al. [71] further develop a system jointly detecting and removing surgical smoke. However, these methods generally simulate surgery smoke for training. Due to the domain gap between the synthetic smoke and the real-world one, they can’t generalize well to real surgery scenes.
@@ 2.2 Supervised Dehazing
Dehazing [5, 29, 19, 11, 63, 30, 47, 67, 37, 64, 72, 45, 20] is of great relevance to surgical desmoking, which aims to recover clean components from hazy ones affected by adverse weather conditions. Li et al. [29] reformulate the atmospheric scattering model and propose AODNet for image dehazing. UHD [63] introduces the infinite approximation of Taylor’s theorem with Laplacian pyramid pattern. DeHamer [19] further combine CNN [48] and vison Transformer [13] together for performance improvement. Compared to the above single-image methods, video ones [30, 47, 67, 67, 37, 64] can leverage temporal clues between consecutive frames for more effective restoration. Li et al. [30] first build an united video detection and dehazing framework that focuses on temporal information fusion. Ren et al. [47] incorporate global semantic priors for smooth transmission map estimation. Recently, MAPNet [64] explores physical haze priors to guide spatial information extraction and proposes a spatial-temporal alignment strategy to guide temporal features aggregation, achieving state-of-the-art performance. However, employing their pre-trained models directly does not produce satisfactory desmoking results due to the discrepancy between atmospheric haze and surgical smoke. Moreover, the lack of real-world paired data severely restricts the possibility of starting training from scratch. In this work, we propose a self-supervised framework for video desmoking (SelfSVD) to address these issues.
@@ 2.3 Unsupervised Desmoking and Dehazing
In contrast to supervised methods, unsupervised ones [73, 57, 24, 43, 50, 8, 53, 31, 32, 70, 65, 33, 15, 66, 16, 9] can be trained without paired supervision. For surgical smoke removal, Cyclic-DesmokeGAN [57] proposes a real-world image desmoking model based on CycleGAN [74]. Desmoke-LAP [43] and DCP-Pixel2Pixel [50] introduce dark channel prior [21] into loss and network design, respectively. MS-CycleGAN [53] adapts a model pre-trained on synthetic data to real-world ones.
@@ 3.1 Taking PS Frame as Supervision
Given a smoky video consisting of $N$ frames $\{\mathbf{S}_{i}\}^{N}_{i=1}$, video desmoking aims to restore the corresponding clean components ${\{\mathbf{\hat{I}}_{i}\}}^{N}_{i=1}$ with temporal clues, i.e.,
@@
$\mathcal{L}_{rec}=\sum_{i=1}^{N}||\mathbf{V}_{i}\odot(\mathbf{\hat{I}}_{i\rightarrow ps}-\mathbf{S}_{ps})||_{1}.$ (5)
$\odot$ is a pixel-wise multiply operation. $\mathbf{V}_{i}$ is a mask that indicates the valid positions of optical flow. The $j$-th element of $\mathbf{V}_{i}$ can be calculated as,
$\mathbf{V}_{i}^{j}=\mathbf{\operatorname{sgn}}\left(\max\left(0,[\mathcal{W}(\mathbf{1},\mathbf{\Psi}_{ps\rightarrow i})]_{j}-\tau\right)\right).$ (6)
$\mathcal{W}$ is warping operation. $\max$ and $\operatorname{sgn}$ are the maximum and sign function, respectively. $\mathbf{1}$ is an all-1 matrix, $[\cdot]_{j}$ denotes the $j$-th element. $\tau$ is set to 0.999.
@@ 3.2 Taking PS Frame as Reference Input
Masking Strategy. In regions with significantly inaccurate optical flow, we prefer to forget the reference information to avoid output contents being inconsistent with smoky inputs. As shown in Fig. 4, we suggest generating a mask $\mathbf{M}_{i}$ to indicate these regions in $\mathbf{F}_{ref\rightarrow i}$, and exclude corresponding features, i.e.,
@@
As shown in Fig. 4, to obtain $\mathbf{M}_{i}$, we first align $\mathbf{S}_{ref}$ to $\mathbf{S}_{i}$, obtaining the warped reference image $\mathbf{S}_{ref\rightarrow i}$. Then, we process $\mathbf{S}_{ref\rightarrow i}$ and $\mathbf{S}_{i}$ with DCP [21] and large-kernel Gaussian blur operations to alleviate the disturbance of surgery smoke. Next, we divide $\mathbf{S}_{ref\rightarrow i}$ and $\mathbf{S}_{i}$ into $P$ non-overlapping patches respectively, i.e., $\{{\mathbf{S}_{ref\rightarrow i}^{p}\}}^{P}_{p=1}$ and $\{{\mathbf{S}_{i}^{p}\}}^{P}_{p=1}$. For each pair $\mathbf{S}_{ref\rightarrow i}^{p}$ and $\mathbf{S}_{i}^{p}$, their structure should be significantly distinct when optical flow is estimated inaccurately. So we adopt structural similarity (SSIM) metric [61] to detect such patch, i.e.,
@@
$m_{i}^{p}$ is the value of $\mathbf{M}_{i}$ in the $p$-th patch. $\epsilon$ is a threshold and is set to 0.92. The patch size is set to $8$. Please see more details about mask in the Suppl.
Figure 5: The illustration of enhancing PS frame ($\mathbf{S}_{ps}$) as supervision. We regard $\mathbf{S}_{ps}$ as a frame with less smoke and feed it into a pre-trained SelfSVD model, generating a cleaner result $\mathbf{S}_{ps}^{\ast}$. Then, $\mathbf{S}_{ps}^{\ast}$ is taken as improved supervision to fine-tune the SelfSVD model by $\mathcal{L}_{rec}$ and $\mathcal{L}_{GAN}$ (i.e., replacing $\mathbf{S}_{ps}$ with $\mathbf{S}_{ps}^{\ast}$ in Eq. 5 and Eq. 11), getting an improved model named SelfSVD.
Regularization Strategy. The areas where optical flow is slightly wrong may lead to potential trivial solutions, as they are not easily detected explicitly. We use $\ell_{1}$ regularization [12, 3] to constraint $\mathbf{F}_{ref\rightarrow i}$ in these areas (indicated by mask $\mathbf{M}_{i}$) to be sparse. The regularization loss $\mathcal{L}_{reg}$ can be written as,
$\mathcal{L}_{reg}=||\mathbf{M}_{i}\odot\mathbf{F}_{ref\rightarrow i}||_{1}.$ (10)
@@ 3.3 Enhancing PS Frame as Supervision
Taking this into account, we improve the training of SelfSVD by enhancing $\mathbf{S}_{ps}$ as better supervision, dubbed SelfSVD, as shown in Fig. 5. In particular, a SelfSVD model is first pre-trained under the supervision of $\mathbf{S}_{ps}$. Then, we regard $\mathbf{S}_{ps}$ as a frame with less smoke and feed it into the pre-trained SelfSVD, getting a cleaner result $\mathbf{S}_{ps}^{\ast}$. Finally, a SelfSVD model is obtained by fine-tuning the pre-trained SelfSVD with $\mathbf{S}_{ps}^{\ast}$ as supervision, and the original SelfSVD model can be detached during testing.
@@ 3.4 Network Architecture and Learning Objective
where $\lambda_{reg}$ and $\lambda_{GAN}$ are set to 0.05 and 1.0, respectively. When training SelfSVD model, $\mathbf{S}_{ps}$ in Eq. 13 is replaced with $\mathbf{S}_{ps}^{*}$, as stated in Sec. 3.3.
4 LSVD Dataset
5 Experiments
5.1 Implementation Details
For optical flow estimation, we adopt a pre-trained PWC-Net [54]. During training, we randomly crop patches and augment them with random flips. The batch size is set to 4 and the patch size is set to $256\times 256$. SelfSVD is trained with ADAM optimizer  [28] with ${\beta}_{1}=0.9$ and ${\beta}_{2}=0.99$ for 100k iterations. Cosine annealing strategy  [38] is employed to decrease the learning rate from $1\times 10^{-4}$ to $1\times 10^{-7}$ steadily. For the training of SelfSVD, we fine-tune the pre-trained SelfSVD model for additional 40k iterations and set the initial learning rate to $5\times 10^{-5}$. All experiments are conducted with PyTorch [44] on a single Nvidia GeForce RTX A6000 GPU.
SelfSVD and SelfSVD keep the same number of residual blocks [22] as BasicVSR++ [7], and the computation costs are generally similar to the video processing methods. Moreover, to make the model inference cost consistent with some single-image processing methods, we present the lightweight versions by reducing the numbers and channels of residual blocks, dubbed SelfSVD-S and SelfSVD-S respectively. Please see more details in the Suppl.
@@ 5.1 Implementation Details
(Ours) SelfSVD 24.58 0.6279 0.4193 4.72 3.86 996 0.18 15.58
@@ 5.3 Comparison with State-of-the-Art Methods
Related learning-based methods are either trained on synthetic data in a supervised manner[29, 59, 11, 63, 19, 20, 6, 7, 64], or trained in a unpaired manner[50, 52, 70, 43]. For fair comparisons, we retrain their models on our LSVD dataset. For training supervised methods, there is no paired ground-truth. Instead, we take the aligned PS frame as the supervision, which is the same as the setting of training our models as stated in Sec. 3.1. Thus, they become self-supervised methods. For unpaired learning methods, we take the original PS frame as their unpaired supervision. Overall, we compare our methods (i.e., SelfSVD-S and SelfSVD-S, SelfSVD, SelfSVD) against 16 related state-of-the-art ones, including 2 unsupervised image processing methods (i.e., PSD [9], DCP [21]), 4 unpaired image ones (i.e., DCP-Pixel2Pixel [50], DistentGAN [52], and Desmoke-LAP [43], RefineNet [70]), 7 self-supervised image ones (i.e., UHD [63], MSDesmoking [59], Wang et al. [60] MSBDN [11], AODNet [29], DADFNet [20], Dehamer [19],) and 3 self-supervised video ones (i.e., BasicVSR [6], BasicVSR++ [7], and MAPNet [64]). We do not compare with unsupervised and unpaired video ones, as few works explore that as far as we know. Beyond conducting experiments on our dataset, we provide more comparisons on a synthetic dataset in the Suppl.
[…]
Quantitative Comparison. Tab. 1 summarizes the quantitative results. First, video processing methods generally perform better than single-image ones, which demonstrates the benefits of utilizing temporal clues for surgical smoke removal. Second, our methods outperform state-of-the-art ones by a large margin. Among all competing methods, BasicVSR++ [7] has achieved the best PSNR score of 23.35dB. Our SelfSVD and SelfSVD achieve PSNR gains of 0.88dB and 1.23dB over BasicVSR++ [7], respectively. Although PSD [70] and DCP [21] get better FADE scores, they generally over-process smoky videos and result in over-saturated colors (as shown in Fig. 6), leading to poor NIQE and PI scores, as well as unsatisfactory visual effects. Third, with lower computation cost and fewer model parameters, the proposed SelfSVD-S and SelfSVD-S still perform well, which further illustrates the effectiveness of our methods.
Qualitative Comparison. Due to space limitation, we only provide qualitative results of some methods with better quantitative scores in Fig. 6. It can be seen that the compared methods often introduce color distortion, over-smoothing, or smoke preservation in their results. Instead, our methods remove more smoke and restore more details that are consistent with PS frames. More qualitative comparison results can be seen in the Suppl.
@@ 6.1 Effect of Taking PS Frame as Supervision
6.2 Effect of Taking PS Frame as Reference Input
Note that PS frame is available during training and testing. We take it as an additional reference (Ref) input. We perform experiments to evaluate its effectiveness by removing the Ref. As shown in Tab. 3, the additional Ref enables 0.67dB PSNR improvements. Besides, Fig. 7 shows that it leads to cleaner smoke removal and finer-scale detail recovery, especially in areas with dense smoke.
We also conduct experiments to validate whether PS frame is a suitable choice to be used as Ref. For comparison, we replace Ref with the first smoky frame, as it generally has less smoke than subsequent smoky ones. As shown in Tab. 3, the replacement leads to 0.28dB PSNR drop, which illustrates the frames before the activations of high-energy devices are preferable to be taken as Ref.
6.3 Effect of Strategies to Avoid Trivial Solutions